OpenStudy (djosh):

Help! Will medal. -- Open Ended Response Given the function f(x) = 4(2)x, Section A is from x = 1 to x = 2 and Section B is from x = 3 to x = 4. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other.

OpenStudy (faiqraees):

Is the function f(x) =4*2x?

OpenStudy (djosh):

f(x) = 4(2)^x, sorry.

OpenStudy (faiqraees):

Start by taking out the first order derivative of f(x)

OpenStudy (anonymous):

Average rate of change is basically asking for the slope of the line through the two points on the graph of the function. \[arc = \frac{ f(b)-f(a) }{ b-a }\] So for your first section, the average rate of change between x = 1 and x = 2 is \[arc = \frac{ f(2)-f(1) }{ 2-1 }\]

OpenStudy (djosh):

Sorry, slope is my weak point. So for section b it would be \[arc = f(4) - f(2) \over 4-1\] based on the previous equation? Sorry if this is wrong, I am just trying to understand and solve the question. @peachpi

OpenStudy (anonymous):

no problem. the second second asks for rate of change between 3 and 4, so they'll be the numbers you use. \[arc = \frac{ f(4)-f(3) }{ 4-3 }\] The numbers you use for each section have to be given to you. You don't use numbers from previous parts. Also, make sure the x values you use in the to numerator and denominator match.

OpenStudy (djosh):

Oh!! I meant to put those numbers, not the 1 & 2! Ah silly me. Typo.

OpenStudy (anonymous):

ok cool. from there you just plug in the x-values into the function to get a final number

OpenStudy (djosh):

I'm in a school with bright lights staring at a computer for a constant 8 hours, so my eyes get all dry and itchy and the headache starts to form, and from there I lose concentration. How do I go about solving the function? Wouldn't the answer just essentially be \[f\]

OpenStudy (anonymous):

Plug in x-values For example for section a \[f(4) = 4(2)^4 = 4(16) = 64\] \[f(3) = 4(2)^3=4(8)=32\] So you have \[arc = \frac{ 64-32 }{ 4-3 }=\frac{32}{1}=32\]

OpenStudy (djosh):

So for section b \[f(2) = 4(2)^2 = 4(4) = 16\] \[f(1) = 4(2)^1 = 4(2) = 8\] So \[arc = \frac{ 16 - 8 }{ 2 - 1} = \frac{ 8 }{ 1 } = 8\]

OpenStudy (anonymous):

yes

OpenStudy (djosh):

THANK YOU SO MUCH FOR YOUR HELP!!

OpenStudy (anonymous):

you're welcome